# Efficient Mixture-of-Expert for Video-based Driver State and Physiological Multi-task Estimation in Conditional Autonomous Driving

### v1

@@ VIII Conclusion
In this study, we introduced the innovative VDMoE model, tailored for comprehensive driver state and physiological measurement estimation, leveraging the rich cues available in facial videos. Our approach uniquely combines the robustness of a MoE framework with the simplicity and efficiency of MLP, and carefully avoids the computational burdens that are often associated with more complex networks like 3DCNNs and Transformers. The VDMoE model is meticulously designed to distill crucial information from facial features. It employs a prior-driven loss function, $\mathcal{L}_{align}$, to align model predictions with known physiological and behavioral patterns, and enhances both accuracy and generalizability. The incorporation of this loss function, alongside the model’s architecture, facilitates a nuanced understanding of both temporal and spatial dynamics in the data, enabling superior performance across tasks of drowsiness, cognitive load, HR, and RR estimation.
Moreover, we presented the MCDD, a comprehensive, large-scale dataset for the monitoring of driver states, designed to explore and estimate the multiple states and the physiological indicators within conditional autonomous driving scenarios, utilizing RGB video data. This dataset was collected based on a driving simulator platform, with strict experiment methodologies employed to accurately reflect the co-occurrence of multi-dimensional cognitive load and drowsiness in real-world conditional autonomous driving environments. We hope that the release of MCDD will benefit future research in this area, particularly in addressing practical real-world concerns. In future work, we will continue to design and collect new multi-task datasets and validate our proposed VDMoE in a real-world conditioned autonomous driving environment.

### v2

@@ VIII Conclusion
In this study, we introduced the innovative VDMoE model, tailored for comprehensive driver state and physiological measurement estimation, leveraging the rich cues available in facial videos. Our approach uniquely combines the robustness of a MoE framework with the simplicity and efficiency of MLP, and carefully avoids the computational burdens that are often associated with more complex networks like 3DCNNs and Transformers. The VDMoE model is meticulously designed to distill crucial information from facial features. It employs a prior-driven loss function, $\mathcal{L}_{align}$, to align model predictions with known physiological and behavioral patterns, and enhances both accuracy and generalizability. The incorporation of this loss function, alongside the model’s architecture, facilitates a nuanced understanding of both temporal and spatial dynamics in the data, enabling superior performance across tasks of drowsiness, cognitive load, HR, and RR estimation. VDMoE is intended as the core component of a real-time DMS in SAE Level-2/3 driving contexts. Given its lightweight architecture and video-based input design, VDMoE can be directly integrated into an in-vehicle system by processing RGB streams from a cabin-facing camera to infer both cognitive load and drowsiness. The model’s multi-task outputs can be fused with vehicle control modules or human-machine interfaces (HMI) to provide adaptive alerts, issue take-over requests, or log physiological states for long-term behavioral analysis. We believe this integration pathway highlights the model’s deployment potential in future intelligent cockpit systems.
Moreover, we presented the MCDD, a comprehensive, large-scale dataset for the monitoring of driver states, designed to explore and estimate the multiple states and the physiological indicators within conditional autonomous driving scenarios, utilizing RGB video data. This dataset was collected based on a driving simulator platform, with strict experiment methodologies employed to accurately reflect the co-occurrence of multi-dimensional cognitive load and drowsiness in real-world conditional autonomous driving environments. We hope that the release of MCDD will benefit future research in this area, particularly in addressing practical real-world concerns. In future work, we will continue to design and collect new multi-task datasets and validate our proposed VDMoE in a real-world conditioned autonomous driving environment. In future work, we will continue to design and collect new multi-task datasets and validate our proposed VDMoE in a real-world conditioned autonomous driving environment, especially in more extreme cases, such as when drivers are experiencing very high cognitive load. Moreover, advancements in aligning drowsiness with cognitive workload distribution by integrating human factor knowledge into data-driven approaches are anticipated. We also plan to explore the integration of VDMoE into real-time embedded platforms and its interaction with existing DMS software pipelines as future work.